Circular No. 6714

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE P. Charles and E. Kuulkers, Oxford University; J. Casares, Instituto

de Astrofísica de Canarias; P. Hakala, University College, London, and Mullard Space Science Laboratory; and P. Muhli, Helsinki University, report: “Optical CCD photometry of V1333 Aql = Aql X-1 with the Nordic Optical Telescope on Aug. 7 shows that it has brightened to V = 17.8 from its quiescent level near mag 19. This brightening occurred during the preceding week, as it was close to quiescence on July 30. Spectroscopy on Aug. 6 with the William Herschel Telescope reveals a blue continuum with superposed Bowen blend, He II 468.6-nm and Balmer emission lines (the Hα equivalent width being ∼ 0.9 nm), confirming the object’s return to activity. An examination of the publicly-available quicklook products of the RXTE ASM shows the presence of weak x-ray emission at a level of 30-40 mCrab beginning between Aug. 1 and 5. This may be a precursor to a major outburst, as occurred six months ago (IAUC 6558). Close monitoring of Aql X-1 is required, especially at different wavelengths.”

COMET C/1997 N1 (TABUR) G. J. Garradd, Loomberah, N.S.W., making CCD observations with

a 0.25-m reflector, reports on the increasing diffuseness of this comet in late July, particularly on July 29, and apparently not just because of the comet’s low altitude in the dawn sky.

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: July 11.94 UT, 12.3 (M. Lehky,

Hradec Kralove, Czech Republic, 0.42-m reflector); 13.97, 12.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 24.19, 12.2 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 29.93, 12.2 (G. Comello, Roden, The Netherlands, 0.30-m reflector); 29.96, 12.31 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); Aug. 3.24, 12.3 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 7.10, 12.2 (J. Bortle, Stormville, NY, 0.41-m reflector).

COMET C/1995 O1 (HALE-BOPP) Total visual magnitude estimates: July 27.89 UT, 4.0 (A. Pearce, Subi-

aco, Western Australia, 20 × 80 binoculars); 30.81, 3.1 (D. A. J. Seargent, The Entrance, N.S.W., naked eye); Aug. 2.35, 4.1 (W. C. de Souza, São Paulo, Brazil, 7 × 50 binoculars).

1997 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden